Find a value of c so that
y equals x squared minus 6 x plus c
y=x2−6x+c has exactly one​ x-intercept.

Answer by mananth(16946)


y=x2−6x+c has exactly one​ x-intercept.
x^2-6x+c =0
The discriminant b^2-4ac =0
here a =1, b=-6 c =c
6^2-4(1)*c=0
36-4c=0
36= 4c
c=9